OpenStudy (anonymous):

Based on the graph of y= x2 - 2x-3, what is the positive x-intercept ?

OpenStudy (anonymous):

writing the equation like this: y = (x-3)(x+1) should help. The x-intercept is when y=0 so just solve this equation: 0 = (x-3)(x+1)

OpenStudy (anonymous):

to find x intercept, set y = 0

OpenStudy (anonymous):

therefore x2 - 2x-3 = 0 and solve for x

OpenStudy (anonymous):

x=-1 or x=3

OpenStudy (anonymous):

thank you